OpenStudy (anonymous):

Identify the inflection point for the function f(x) = 8sin x + 2x^2.

OpenStudy (anonymous):

Take second derivative and set it equals to 0, that's your inflection point

OpenStudy (bahrom7893):

No whenever the signs change

OpenStudy (bahrom7893):

Like that too and also signs have to change

myininaya (myininaya):

right it usually happens when f''=0 but we can always choose test numbers before and after each possible inflection point to see if the concavity switches and if it does then it is an inflection point

myininaya (myininaya):

so what is f' curiousmath?

myininaya (myininaya):

(8sinx)'=8(sinx)'=8*? (2x^2)'=?

myininaya (myininaya):

fill in the question marks

OpenStudy (anonymous):

8cosx and 4x

OpenStudy (anonymous):

am i wrong ?

myininaya (myininaya):

good!

myininaya (myininaya):

ok so now f'(x)=8cosx+4x now we need f'' f''(x)=-8cosx+4

myininaya (myininaya):

now set =0 and solve oh yeah do you have any restrictions on the domain

OpenStudy (anonymous):

okay so now we have to make it equal to 0. and find the three points right?

myininaya (myininaya):

there is infinitly many possible inflection points

myininaya (myininaya):

so we have -8cosx+4=0 so -8cosx=-4 cosx=4/8 cosx=1/2 now what x's can you think of that will make this true?

OpenStudy (anonymous):

Isn't f''(x) = -8sinx + 4

myininaya (myininaya):

yes you are right

OpenStudy (anonymous):

I feel like ripping my hair out. omg I have no idea!

myininaya (myininaya):

-8sinx+4=0 -8sinx=-4 sinx=4/8 sinx=1/2

myininaya (myininaya):

what about when x=pi/6 and 5pi/6 if these are inflection points then so are x=pi/6+2npi and x=5pi/6+2npi for n=...,-3,-2,-1,0,1,2,3,....

myininaya (myininaya):

and if you plug in number before and after pi/6 you see the concavity switches and if you plug in number before and after 5pi/6 you see the concavity switches so (pi/6+2npi,f(pi/6+2npi)) and (5pi/6+2npi, f(5pi/6+2npi)) are inflection points

myininaya (myininaya):

when you pick your test numbers make sure its in the intervals between the possible inflection points and the areas between the possible inflection points and infinity